### v1

# DART: A Diffusion-Based Autoregressive Motion Model for Real-Time Text-Driven Motion Control

Abstract
Text-conditioned human motion generation, which allows for user interaction through natural language, has become increasingly popular. Existing methods typically generate short, isolated motions based on a single input sentence. However, human motions are continuous and can extend over long periods, carrying rich semantics. Creating long, complex motions that precisely respond to streams of text descriptions, particularly in an online and real-time setting, remains a significant challenge. Furthermore, incorporating spatial constraints into text-conditioned motion generation presents additional challenges, as it requires aligning the motion semantics specified by text descriptions with geometric information, such as goal locations and 3D scene geometry. To address these limitations, we propose DART, a Diffusion-based Autoregressive motion primitive model for Real-time Text-driven motion control. Our model, DART, effectively learns a compact motion primitive space jointly conditioned on motion history and text inputs using latent diffusion models. By autoregressively generating motion primitives based on the preceding history and current text input, DART enables real-time, sequential motion generation driven by natural language descriptions. Additionally, the learned motion primitive space allows for precise spatial motion control, which we formulate either as a latent noise optimization problem or as a Markov decision process addressed through reinforcement learning. We present effective algorithms for both approaches, demonstrating our model’s versatility and superior performance in various motion synthesis tasks. Experiments show our method outperforms existing baselines in motion realism, efficiency, and controllability.
1 Introduction
Text-conditioned human motion generation has gained increasing popularity in recent years for flexible user interaction via natural languages. Existing text-conditioned motion models  (Tevet et al., 2023; Guo et al., 2023; Zhang et al., 2023a; Guo et al., 2022; Jiang et al., 2024a) primarily focus on generating standalone short motions from a single descriptive sentence, thereby overlooking the inherently sequential nature of human motion. These methods fail to generate long and complex motions consisting of multiple segments with distinct conditions of action descriptions. FlowMDM (Barquero et al., 2024) is the state-of-the-art temporal motion composition method, capable of generating complex, continuous motions while providing control over the desired actions and their durations. However, FlowMDM is an offline method that requires prior knowledge of the entire action timeline and has a slow generation speed, making it unsuitable for online and real-time applications.
In addition to text-based semantic control, generating human motion within spatial constraints and achieving specific goals—such as reaching a keyframe body pose, following the joint trajectory, or interacting with objects-offers broad applications but introduces additional complex challenges. Recent works (Shafir et al., 2024; Karunratanakul et al., 2024; Xie et al., 2024) have sought to integrate text-conditioned motion models with spatial control capabilities. However, they still face challenges in simultaneously harmonizing spatial control, motion quality, and semantic alignment with text. Moreover, these approaches are typically restricted to controlling isolated short motions in an offline setting. Conversely, interactive character control (Kovar et al., 2008; Holden et al., 2015; Ling et al., 2020; Peng et al., 2022) has been a longstanding focus in computer graphics, with a primary emphasis on achieving motion realism and responsiveness to interactive control signals. However, most of these works lack support for text-conditioned semantic control and are limited by being trained on small, curated datasets. Incorporating text-conditioned motion generation could provide a novel, efficient language interface for animators and everyday users to control the characters, reducing the effort required when specifying spatial control signals is challenging or tedious.
To address these limitations, we propose DART, a diffusion-based autoregressive motion primitive model for real-time text-conditioned motion composition. Moreover, the compact and expressive motion space of DART provides a foundation for integrating precise spatial control through latent space optimization or reinforcement learning-based algorithms. DART features three key components.
First, DART represents human motion as a collection of motion primitives (Zhang & Tang, 2022), which are autoregressive representations consisting of overlapping short motion segments tailored for online generation and control. These short primitives also provide a clearer alignment with atomic action semantics compared to longer sequences, enabling effective learning of a text-conditioned motion space. By focusing on shorter primitives, DART avoids the complexities and extensive data demands of modeling entire motion sequences, allowing for high-quality motion generation with only a few diffusion steps.
Next, DART learns a text-conditioned motion primitive space from large-scale data using a latent diffusion architecture, which contains a denoiser network for generating motions in a compact latent motion space, and a decoder network for reconstructing the detailed motions. Leveraging this learned latent motion space, DART employs an autoregressive rollout to synthesize motion sequences from real-time text inputs, enabling the efficient generation of motions of arbitrary length. Compared with the offline temporal motion composition method FlowMDM (Barquero et al., 2024), DART provides real-time response and ten times faster generation speed.
Lastly, the versatile and powerful motion space learned in DART enables precise spatial control through latent space optimization or RL-based policy learning. We present effective learning and optimization algorithms that explore the learned DART motion space to obtain motion sequences that precisely follow both textual and spatial constraints. We evaluate DART across various motion synthesis tasks, including generating long, continuous sequences from sequential text prompts, in-between motion generation, scene-conditioned motion, and goal-reaching synthesis. The experimental results show that DART is a simple, unified and highly effective motion model, consistently outperforming or matching the performance of baselines.
2 Related Works
Conditional Motion Generation. Generating realistic and diverse human motions is a long-standing challenge in computer vision and graphics. Apart from generating highly realistic human motions (Kovar et al., 2008; Holden et al., 2020; Büttner & Clavet, 2015; Zinno, 2019), conditional generation is another important factor that aligns motion generation with human intentions and various application constraints. Text-conditioned motion generation (Tevet et al., 2023; Zhang et al., 2022; Petrovich et al., 2022; Guo et al., 2022; Jiang et al., 2024a; Zhang et al., 2023a) has become increasingly popular since it allows users to modulate motion generation with flexible natural languages. Audio and speech-driven motion synthesis methods (Alexanderson et al., 2023; Tseng et al., 2022; Siyao et al., 2022; Ao et al., 2022; 2023) have also made significant progress recently. Moreover, there exist many applications that require spatial awareness and precise control in motion generation, such as interactive character control (Ling et al., 2020; Peng et al., 2022; Starke et al., 2022; Luo et al., 2023; Starke et al., 2024), human-scene interactions(Hassan et al., 2021; Starke et al., 2019; Zhao et al., 2022; 2023; Li et al., 2024; Xu et al., 2023; Jiang et al., 2024b; Wang et al., 2024; Liu et al., 2023; Li et al., 2023a), and human-human(noid) interactions (Liang et al., 2024; Zhang et al., 2023c; Christen et al., 2023; Cheng et al., 2024; Shan et al., 2024). Synthesizing high-quality motions with precise spatial control remains challenging, and DART is a step toward a general and efficient motion model that supports precise control tasks.
Diffusion Generative Models. Denoising Diffusion Models (Ho et al., 2020; Song et al., 2020a; b) are generative models that learn to predict clean data samples by gradually annealing the noise from a standard Gaussian sample. Diffusion models have achieved unprecedented performances in many generation tasks including images, videos, and 3D human motions (Tevet et al., 2023; Rombach et al., 2022; Ho et al., 2022). Diffusion models can accept flexible conditions to modulate the generation, such as text prompts, images, audio, and 3D objects (Rombach et al., 2022; Tevet et al., 2023; Alexanderson et al., 2023; Tseng et al., 2022; Zhang et al., 2023b; Xu et al., 2023; Li et al., 2023b). Most existing diffusion-based motion generation methods focus on offline generations (Tevet et al., 2023; Barquero et al., 2024; Chen et al., 2023; Karunratanakul et al., 2024; Cohan et al., 2024; Dai et al., 2024) producing short, isolated motion sequences at once while neglecting the autoregressive nature of human motions. Among these methods, DNO (Karunratanakul et al., 2024) is closely related to our optimization-based control approach, as both use diffusion noises as the latent space for editing and control. However, the key distinction lies in our latent motion primitive-based diffusion model, which, unlike DNO’s diffusion model trained with full motion sequences in raw space, achieves superior performance in harmonizing spatial control with text semantic alignment during experiments. There are also works that incorporate history conditions in diffusion-based motion prediction. Recent works (Xu et al., 2023; Jiang et al., 2024b) use history-conditioned diffusion models to autoregressive generate human-scene interaction sequences. Van Wouwe et al. (2024) uses autoregressive diffusion for real-time human motion reconstruction from IMU sensors. Shi et al. (2024) and Chen et al. (2024) adapt diffusion models for real-time character motion generation and control. While Shi et al. (2024) learns character control policies using diffusion noises as the action space, akin to our reinforcement-learning-based control, their method focuses on single-frame autoregressive generation and lacks support for text conditions, which offer a compact and powerful interface for users to control character behaviors. In contrast, DART is an efficient and general motion model that scales effectively to large motion datasets with text annotations (Mahmood et al., 2019; Punnakkal et al., 2021). DART supports natural language interfacing and provides a versatile foundation for various motion generation tasks with spatial control.
3 Method
3.1 Preliminaries
Problem Definition. We focus on the task of text-conditioned online motion generation with spatial control. Given an $H$ frame seed motion $\mathbf{H}_{seed}=[\mathbf{h}^{1},...,\mathbf{h}^{H}]$, a sequence of $N$ text prompts $C=[c^{1},...,c^{N}]$, and spatial goals $g$, the objective is to autoregressively generate continuous and realistic human motion sequences $\mathbf{M}=[\mathbf{H}_{seed},\mathbf{X}^{1},...,\mathbf{X}^{N}]$, where each motion segment $\mathbf{X}^{i}$ matches the semantics of the corresponding text prompt $c^{i}$ and satisfies the spatial goal constraints $g$. This task imposes challenges in high-level action semantic control, precise spatial control, and smooth temporal transition in motion generation.
Autoregressive Motion Primitive Representation. We model long-term human motions as the sequential composition of motion primitives (Zhang & Tang, 2022) with overlaps for efficient generative learning and online inference. In contrast to directly modeling long motion sequences, the primitive representation decomposes globally complex sequences into short and simple primitives, resulting in a more tractable data distribution for generative learning. This enables DART to learn high-quality primitive generation using as few as 10 diffusion steps. Moreover, the autoregressive and simple nature of the primitive representation makes it inherently suitable for fast online generations. Furthermore, motion primitives convey more interpretable semantics than individual frames, enhancing the learning of text-conditioned motion space.
Each motion primitive $\mathbf{P}^{i}=[\mathbf{H}^{i},\mathbf{X}^{i}]$ is a short motion clip containing $H$ frames of history motion $\mathbf{H}^{i}=[\mathbf{h}^{i,1},...,\mathbf{h}^{i,H}]$ that overlap with the previous motion primitive, and $F$ frames of future motion $\mathbf{X}^{i}=[\mathbf{x}^{i,1},...,\mathbf{x}^{i,F}]$. The history motion of the i-th motion primitive $\mathbf{H}^{i}$ consists of the last $H$ frames of the previous motion primitive $\mathbf{X}^{i-1,F-H+1:F}$. Therefore, infinitely long motions can be represented as the rollout of such overlapping motion primitives as $\mathbf{M}=[\mathbf{H}_{seed},\mathbf{X}^{1},...,\mathbf{X}^{N}]$.
We represent each motion frame using an overparameterized representation based on the SMPL-X (Pavlakos et al., 2019) parametric human body model. The overparameterized representation is a $D=276$ dimensional vector including the body root translation $\mathbf{t}$, root orientation $\mathbf{R}$, local joint rotations $\mathbf{\theta}$, joint locations $\mathbf{J}$, and the temporal difference features of locations and rotations. All the frames within one motion primitive are canonicalized in a local coordinate frame centered at the first-frame body of the primitive. We use history length $H=2$ and future length $F=8$ in our experiments. Further details of the primitive representation are attached in the Appendix A.
3.2 DART: A Diffusion-Based Autoregressive Motion Primitive Model
Figure 1: Architecture illustration of DART. The encoder network compresses the future frames $\mathbf{X}=[\mathbf{x}^{1},...,\mathbf{x}^{F}]$ into a latent variable, conditioned on the history frames $\mathbf{H}=[\mathbf{h}^{1},...,\mathbf{h}^{H}]$. The decoder network reconstructs the future frames conditioned on the history frames and latent sample $z$. The denoiser network predicts the clean latent sample $\hat{z}_{0}$ conditioned on the noising step, text prompt, history frames, and noised latent sample $z_{t}$. During the denoiser training, both the encoder and decoder networks remain frozen.
We propose a novel latent diffusion model (Rombach et al., 2022; Chen et al., 2023) designed for motion primitives. This model can autoregressively generate motions with smooth transitions, conditioned on text prompts and motion history. The proposed model contains a variational autoencoder (VAE) (Kingma & Welling, 2022) that compresses the motion primitives into a compact latent space, and a latent denoising diffusion model that predicts clean latent variables from noises, conditioned on text prompts and motion history.
Latent Motion Primitive Space. We introduce a motion primitive VAE that compresses motion primitives into a compact latent space, upon which we train our latent diffusion models, rather than using the raw motion space (Rombach et al., 2022; Chen et al., 2023). This design is inspired by the observations that raw motion data from motion capture datasets (Mahmood et al., 2019) often contain various levels of high-frequency noises, and training diffusion models on raw motion space leads to results inheriting such noise artifacts. This is evidenced by the significantly higher jittering in generated motions of the ablative model without VAE in Appendix. E.3. The compression achieved through the motion primitive VAE significantly mitigates the impact of high-frequency noise. The resulting latent representation is not only more compact but also more computationally efficient than the raw motion data, thereby enhancing the efficiency of our generative model and improving the control capabilities within the latent space.
Our motion primitive VAE employs a transformer-based architecture based on Chen et al. (2023), comprising an encoder $\mathcal{E}$ and decoder $\mathcal{D}$, as shown in Fig. 1. The encoder takes as input the history motion frames $\mathbf{H}$ and future motion frames $\mathbf{X}$ as well as the learnable distribution tokens $T_{\mu}$ and $T_{\sigma}$, which are responsible for predicting the latent distribution. The latent sample $z$ is obtained from the predicted distribution via reparameterization (Kingma & Welling, 2022). The decoder then predicts the future frames $\mathbf{\hat{X}}$ from zero tokens conditioned on the latent sample $z$ and the history frames $\mathbf{H}$. The motion primitive VAE is trained with the future frame reconstruction loss $L_{rec}$, auxiliary losses $L_{aux}$ that penalize unnatural motion reconstruction, and a small Kullback-Leibler (KL) regularization term $L_{KL}$. We refer to Appendix C for further details about the motion primitive VAE.
Latent Motion Primitive Diffusion. Building on the motion representation provided by our trained motion primitive encoder $\mathcal{E}$, we transform the motion primitives from the raw feature space to a more refined latent space that is more suitable for generative learning with diffusion models. This latent motion primitive space is then used to train a diffusion denoiser for generating motion primitives conditioned on text prompts. The denoiser architecture is shown in Fig. 1. Given the noised latent variable $z_{t}$, the diffusion step $t$, the history motion $\mathbf{H}$, and a text prompt $c$, we train a transformer-based latent denoiser model $\mathcal{G}$ to predict the cleaned motion primitive latent variable $\hat{z}_{0}=\mathcal{G}(z_{t},t,\mathbf{H},c)$. The diffusion step is embedded using a small MLP, while the text prompt is encoded using the CLIP (Radford et al., 2021) text encoder. The text prompt is randomly masked out by a probability of 0.1 during training to enable classifier-free guidance (Ho & Salimans, 2022) during generation. During training, the motion primitive latent variable $z_{0}=\mathcal{E}(\mathbf{H},\mathbf{X})$ can be efficiently obtained with a single inference of the fixed encoder $\mathcal{E}$. The cleaned latent variable $\hat{z}_{0}$ can be converted back to the future frames using the fixed decoder $\mathcal{D}$: $\hat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\hat{z}_{0})$. The latent denoiser $\mathcal{G}$ is trained with the simple objective (Ho et al., 2020). Additionally, we use the reconstruction loss $L_{rec}$ and auxiliary losses $L_{aux}$ on the decoded future frames $\hat{\mathbf{X}}$, the same as training the motion primitive VAE. Notably, we use DDPM (Ho et al., 2020) with only 10 diffusion steps during both the training and inference of the denoiser model. This is sufficient for generating realistic samples due to the simplicity of the short motion primitives and enables highly efficient online generation. Moreover, we use the scheduled training (Ling et al., 2020; Bengio et al., 2015; Rempe et al., 2021) to progressively introduce the test-time distribution of the history motion $\mathbf{H}$, which improves the stability of long sequence online generation and the text prompt controllability for unseen poses. We refer to Appendix D for the details.
With the trained motion primitive decoder $\mathcal{D}$, latent denoiser $\mathcal{G}$ and a diffusion sampler $\mathcal{S}$ such as Ho et al. (2020) and Song et al. (2022), we can autoregressively generate motion sequences given the history motion seed $\mathbf{H}_{seed}$ and the online sequence of text prompts $C$, as shown in Alg. 1. During sampling, we use classifier-free guidance (Ho & Salimans, 2022) on the text condition with a guidance scale $w$:
$\mathcal{G}_{w}(z_{t},t,\mathbf{H},c)=\mathcal{G}(z_{t},t,\mathbf{H},\emptyset)+w\cdot(\mathcal{G}(z_{t},t,\mathbf{H},c)-\mathcal{G}(z_{t},t,\mathbf{H},\emptyset))$ (1)
Using the rollout algorithm, DART generates over 300 frames per second using a single RTX 4090 GPU, enabling real-time applications and online reinforcement-learning control as in Sec. 3.3.
Algorithm 1 Autoregressive rollout generation using latent motion primitive model
Input: primitive decoder $\mathcal{D}$, latent variable denoiser $\mathcal{G}$, history motion seed $\mathbf{H}_{seed}$, text prompts $C=[c^{1},...,c^{N}]$, total diffusion steps $T$, classifier-free guidance scale $w$, diffusion sampler $\mathcal{S}$.
Optional Input: Latent noises $Z_{T}=[z_{T}^{1},...,z_{T}^{N}]$
Output: motion sequence $\mathbf{M}$
$\mathbf{H}\leftarrow\mathbf{H}_{seed}$
$\mathbf{M}\leftarrow\mathbf{H}_{seed}$
for $i\leftarrow 1$ to $N$ do $\triangleright$ number of rollouts
     sample noise $z_{T}^{i}$ from $\mathcal{N}(0,1)$ if not inputted
     $\hat{z}_{0}^{i}\leftarrow\mathcal{S}(\mathcal{G},z_{T}^{i},T,\mathbf{H},c^{i},w$) $\triangleright$ diffusion sample loop with classifier-free guidance
     $\hat{\mathbf{X}}\leftarrow\mathcal{D}(\mathbf{H},\hat{z}_{0}^{i})$
     $\mathbf{M}\leftarrow\textsc{concat}(\mathbf{M},\hat{\mathbf{X}})$triangleright concatenate future frames to generated sequence
     $\mathbf{H}\leftarrow\textsc{canonicalize}(\hat{\mathbf{X}}^{F-H+1:F})$triangleright update the rollout history with last H generated frames
end for
return $\mathbf{M}$
3.3 Controllable Motion Synthesis via DART
Text-conditioned motion generation offers a user-friendly interface for controlling motions through natural language. However, relying solely on text limits precise spatial control, such as walking to a specific location or sitting in a designated spot. Therefore, it is necessary to incorporate motion control mechanisms to achieve precise spatial goals, including reaching a keyframe body pose, following joint trajectories, and interacting with scene objects.
We formulate the motion control task as generating the motion sequence $\mathbf{M}$ that minimizes its distance to a given spatial goal $g$ under a criterion function $\mathcal{F}(\cdot,\cdot)$ and the regularization from the scene and physical constraints $cons(\cdot)$:
$\mathbf{M}^{*}={\arg\!\min}_{\mathbf{M}}\mathcal{F}(\Pi(\mathbf{M}),g)+cons(\mathbf{M})$ (2)
where $g$ is the task-dependent spatial goal, e.g., $g$ is a keyframe body for the motion in-between task or the goal location for the navigation task. $\Pi(\cdot)$ is the goal-dependent projection function that maps motion sequence to the goal observation space, $cons(\cdot)$ denotes physical constraints such as avoiding collisions into scenes.
Directly solving the motion control task in the raw motion space often results in unrealistic motions since most samples in the raw motion space do not represent plausible motions. To improve the generated motion quality, many previous methods teacle such motion control tasks in a latent motion space, where samples can be mapped to plausible motions (Karunratanakul et al., 2024; Ling et al., 2020; Peng et al., 2022; Holden et al., 2015). DART offers a powerful text-conditioned latent motion space for such latent space control. DART learns a powerful generative model that can generate diverse and realistic motions from standard Gaussian samples $Z_{T}$. Using the deterministic DDIM (Song et al., 2022) sampler, we adapt DART sampling to function as a deterministic mapping from latent noises $Z_{T}$ to plausible motions. Using the latent noises $Z_{T}$ as the latent space, we transform the motion control task in Eq. 2 to latent space control as follows: Given the initial motion history $\mathbf{H}_{seed}$, a sequence of text prompts $C$, and a deterministic diffusion sampler $\mathcal{S}$, the rollout function in Alg. 1 can deterministically map a list of motion primitive latent noises $Z_{T}=[z_{T}^{1},..,z_{T}^{N}]$ to a motion sequence $\mathbf{M}$ by :
$\mathbf{M}=\textsc{rollout}(Z_{T},\mathcal{D},\mathcal{G},\mathbf{H}_{seed},C,T,w,\mathcal{S})$ (3)
We abbreviate it as $\textsc{rollout}(Z_{T},...)$ for short since $Z_{T}$ is the only variable in latent control. The minimization objective is converted as :
$Z_{T}^{*}={\arg\!\min}_{Z_{T}}\mathcal{F}(\Pi(\textsc{rollout}(Z_{T},...),g)+cons(\textsc{rollout}(Z_{T},...))$ (4)
Note that we do not use DDIM to skip diffusion steps at sampling, which we observe to cause artifacts in generated motion. We then propose two solutions to this latent space motion control problem, one is to directly optimize the latent noises using gradient descent, and the other is to model the control task as a Markov process and use reinforcement learning to learn control policies.
Motion Control via Latent Diffusion Noise Optimization. One straightforward solution to this minimization problem (Eq. 4) is to directly optimize the latent noises $Z_{T}$ given the criterion function using gradient descent methods. The latent noise optimization is illustrated in Alg. 2.
Algorithm 2 Latent noises optimization
Input: Latent noises $Z_{T}=[z_{T}^{1},...,z_{T}^{N}]$, Optimizer $\mathcal{O}$, learning rate $\eta$, criterion $\mathcal{L}$, goal $g$ (for brevity, we do not reiterate the inputs of the rollout function defined in Alg. 1)
Output: motion sequence $\mathbf{M}$
for $i\leftarrow 1$ to optimization steps do
     $\mathbf{M}\leftarrow\textsc{rollout}(Z_{T},...)$
     $\nabla\leftarrow\nabla_{Z_{T}}\mathcal{L}(\mathbf{M})$
     $Z_{T}\leftarrow\mathcal{O}(Z_{T},\nabla/\|\nabla\|,\eta)$triangleright update using normalized gradient
end for
return $\mathbf{M}\leftarrow\textsc{rollout}(Z_{T})$
This optimization-based control framework is general and applicable for various spatial control tasks. We instantiate the latent noise optimization method in two example control scenarios: in-between motion generation and human-scene interaction generation. First, we addresss the motion in-between task that aims to generate the motion frames transition between given history and goal keyframes $g$ that is $f$ frames away conditioned on the text prompt $c$. We use the distance between the $f$-th frame of the generated motion and the goal keyframe as the optimization objective.
Second, we show that physical and scene constraints $cons(\cdot)$ can be incorporated to synthesize human motions in a contextual environment. Given the text prompts $C$ and spatial goal $g$ of the interaction anchor joint locations, e.g., locations of the pelvis when sitting, the objective is to generate motions that not only perform the desired interaction but also achieve the goal joint positions while adhering to scene constraints. We consider two scene constraints penalizing the human-scene collision and foot-floating artifacts, as detailed in Appendix. F.
Motion Control via Reinforcement Learning. Although the proposed latent noise optimization is effective for general control tasks, the optimization can be computationally expensive. To address this, the autoregressive primitive-based motion representation of DART allows for another efficient control mechanism using reinforcement learning (RL) (Sutton & Barto, 1998). We model the latent motion control as a Markov decision process with a latent action space and use RL to learn policy models to achieve the goals. We model digital humans as agents to interact with an environment according to a policy to maximize the expected discounted return. At each time step $i$, the agent observes the state $\mathbf{s}^{i}$ of the system, samples an action $\mathbf{a}^{i}$ from the learned policy, with the system transitioning to the next state $\mathbf{s}^{i+1}$ due to the performed action $\mathbf{a}^{i}$, and receives a reward $r^{i}=r(\mathbf{s}^{i},\mathbf{a}^{i},\mathbf{s}^{i+1})$.
Figure 2: Architecture of the reinforcement learning-based control policy. The frozen DART diffusion denoiser and decoder models transform the latent actions into motion frames. The last predicted frames are canonicalized and provided to the policy model as the next step history condition.
Our latent motion primitive model naturally fits in the Markov decision process due to its autoregressive nature. We use the latent noises $z_{T}$ as the latent action space and train goal-conditioned policy models as controllers. The policy architecture is shown in Fig. 2. Our policy model employs the actor-critic (Sutton & Barto, 1998) architecture and is trained with the PPO (Schulman et al., 2017) algorithm. The state $\mathbf{s}^{i}$ includes the history motion observation $\mathbf{H}^{i}$, goal observation $g^{i}$, scene observation $s^{i}$, and the CLIP embedding of the text prompt $c^{i}$. The policy model takes in [$\mathbf{H}^{i}$, $g^{i}$, $s^{i}$, $c^{i}$] to predict the latent noise $z_{T}^{i}$ as the action. The latent noise $z_{T}^{i}$ is mapped to the future motion frames $\mathbf{X}^{i}$ using the frozen latent denoiser $\mathcal{G}$ and motion primitive decoder $\mathcal{D}$. The new history motion is extracted from the last $H$ predicted frames and fed to the policy network in the next step. We reformulate the minimization problem in Eq. 4 as reward maximization to train the policy.
We instantiate the reinforcement learning control with the text-conditioned goal-reaching task. Given a text prompt $c$ and a 2D goal location $g$, we aim to control the human to reach the goal location using the action specified by the text. The goal location is converted to a local observation consisting of its distance to the human pelvis, and its local direction in the human-centric coordinates frame. We consider a simple flat scene and the scene observation is the relative floor height to the human pelvis. The policy is trained with distance rewards encouraging the human pelvis to reach the goal location and scene constraint rewards penalizing foot skating and floor penetration. Further details can be found in appendix G. With the trained control policies, we can efficiently control a human to reach dynamic goals using specified skills like walking or hopping.
4 Experiments
We provide extensive experiments showing how DART can serve as a general model for text-conditioned temporal motion composition (4.1) and various motion generation tasks requiring precise spatial control via latent noise optimization (4.2) and reinforcement learning policy (4.3). The qualitative results and comparisons are attached in the supplementary video.
Our DART is trained on motion-text data from the BABEL (Punnakkal et al., 2021) dataset in our experiments if not otherwise stated. BABEL contains motion capture sequences with frame-aligned text labels that annotate the fine-grained semantics of actions. Fine-grained text labels in BABEL allow models to learn precise human action controls and natural transitions among actions. However, DART can also learn using motions with coarse sequence-level labels such as the HML3D (Guo et al., 2022) dataset, as in the optimization-based motion in-between experiments in Sec. 4.2.
4.1 Text-Conditioned Temporal Motion Composition
Text-conditioned temporal motion composition aims to generate realistic motion sequences that faithfully align with a list of action segments, each defined by a specific text prompt and duration. We evaluate the motion composition task on the BABEL(Punnakkal et al., 2021) dataset consisting of motion capture sequences with human-annotated per-frame action descriptions, which facilitate the evaluation of precise action controls and natural transitions in motion composition. Since BABEL does not release the test set, we compare our DART with baseline methods on the BABEL validation set. We extract the list of action segments described by tuples of text prompts and durations from each data sequence and feed the action lists as conditions for motion composition.
We evaluate the generation results using metrics proposed in Guo et al. (2022) and Barquero et al. (2024). For each action segment, we evaluate the similarity between generation and dataset (FID), motion-text semantic alignment (R-prec, MM-DIST), and generation diversity (DIV). To evaluate smooth transitions between two segments, we measure the jerk (the derivative of acceleration) of the 30-frame transition clip centered at the splitting point of two action segments, reporting the peak jerk (PJ) and Area Under the Jerk (AUJ). Moreover, we profile all methods in a benchmark of generating one 5000-frame-long sequence and report the generation speed, the latency of getting the first generated frames, and memory usage. We also conduct binary user studies to evaluate the motion realism and motion-text semantic alignment of generation results, during which the participants are given generation results from two different methods and are asked to select the generation that is perceptually more realistic or better aligns with the action text stream in subtitles. We compare DART with baselines inlcuding TEACH (Athanasiou et al., 2022), DoubleTake (Shafir et al., 2024), a history-conditioned modification of T2M-GPT (Zhang et al., 2023a)(denoted as T2M-GPT*), and the state-of-the-art offline motion composition method FlowMDM (Barquero et al., 2024).
We present the quantitative results in Tab. 1 and Tab. 2. DART achieves the best FID in both the segment and transition evaluation, indicating the highest similarity to the dataset and best motion realism. DART also displays second-best jerk metrics indicating smooth action transitions. We observe that DART performs slightly worse than FlowMDM in motion-text semantic alignment (R-prec and MM-Dist) because of the online generation nature of DART. Natural action transitions require time to transit to the new action after receiving the new action prompt, leading to a delay in the emergence of the new action semantics. For instance, a human cannot immediately transition from kicking a leg in the air to stepping backward without first recovering to a standing pose. This inherent delay in transitions results in a motion embedding shift that impacts the R-prec metric of DART. In contrast, the offline baseline FlowMDM generates the entire sequence as a whole and requires oracle information of the full timeline of action segments to modulate compatibility between subsequent segments. The slight but natural action transition delay of DART is perceived as natural by humans, as shown in the user study results in Tab. 2 where DART is preferred over all the baselines, including FlowMDM, for both motion realism and motion-text semantic alignment in human evaluations.
DART requires significantly less memory than the offline baseline FlowMDM and achieves approximately 10x faster generation, with a frame rate exceeding 300 and a latency of 0.02s, enabling real-time text-conditioned motion composition (see supplementary video). We refer to Appendix E for details of experiments, user studies, and ablation studies about model architecture and hyperparameters, and refer to the supplementary video for qualitative comparisons.
Table 1: Quantitative evaluation results on text-conditioned temporal motion composition. The first row includes the metrics of the dataset for reference. Symbol $\rightarrow$’ denotes that closer to the dataset reference is better and ‘$\pm$’ indicates the 95% confidence interval. Bold and blue texts indicate the best and second best results excluding the dataset, respectively.
Segment Transition Profiling
[…]
Dataset 0.00$\pm$0.00 0.72$\pm$0.00 8.42$\pm$0.15 3.36$\pm$0.00 0.00$\pm$0.00 6.20$\pm$0.06 0.02$\pm$0.00 0.00$\pm$0.00
TEACH 17.58$\pm$0.04 0.66$\pm$0.00 10.02$\pm$0.06 5.86$\pm$0.00 3.89$\pm$0.05 5.44$\pm$0.07 1.39$\pm$0.01 5.86$\pm$0.02 3880$\pm$144 0.05$\pm$0.00 2251
DoubleTake 7.92$\pm$0.13 0.60$\pm$0.01 8.29$\pm$0.16 5.59$\pm$0.01 3.56$\pm$0.05 6.08$\pm$0.06 0.32$\pm$0.00 1.23$\pm$0.01 85$\pm$1 59.11$\pm$0.76 1474
T2M-GPT* 7.71$\pm$0.55 0.49$\pm$0.01 8.89$\pm$0.21 6.69$\pm$0.08 2.53$\pm$0.04 6.61$\pm$0.02 1.44$\pm$0.03 4.10$\pm$0.09 885$\pm$12 0.23 $\pm$0.00 2172
FlowMDM 5.81$\pm$0.10 0.67$\pm$0.00 8.90$\pm$0.06 5.08$\pm$0.02 2.39$\pm$0.01 6.63$\pm$0.08 0.04$\pm$0.00 0.11$\pm$0.00 31$\pm$0 161.29$\pm$0.24 11892
Ours 3.79$\pm$0.06 0.62$\pm$0.01 8.05$\pm$0.10 5.27$\pm$0.01 1.86$\pm$0.05 6.70$\pm$0.03 0.06$\pm$0.00 0.21$\pm$0.00 334 $\pm$2 0.02$\pm$0.00 2394
Table 2: Binary user study results comparing our method against baselines in generation realism and motion-text semantic alignment on text-conditioned temporal motion composition. We report the percentage of each method being voted better than the other (Ours vs. Baselines).
Realism (%) Semantic (%)
Ours vs. TEACH 66.7 vs. 33.3 66.0 vs. 34.0
Ours vs. DoubleTake 66.4 vs. 33.6 66.1 vs. 33.9
Ours vs. T2M-GPT* 61.3 vs. 38.7 66.7 vs. 33.3
Ours vs. FlowMDM 53.3 vs. 46.7 51.3 vs. 48.7
4.2 Latent Diffusion Noise Optimization-Based Control Using DART
Text-conditioned motion in-between.
Motion in-betweening aims to generate realistic motion frames that smoothly transition between a pair of history and goal keyframes. We consider a text-conditioned variant where an additional text prompt is inputted to specify the action semantics of the frames in between. We compare our method with DNO (Karunratanakul et al., 2024) and OmniControl (Xie et al., 2024). For a fair comparison, we train DART on the HML3D dataset same as the baselines. We evaluate using test sequences covering diverse actions, with the sequence lengths ranging from 2 to 4 seconds. The quantitative evaluations are shown in Tab. 3. We report the $L_{2}$ norm errors between the generated motion and the history motion and goal keyframe. We also evaluate the motion realism with the skate and jerk metrics. The skate metric (Ling et al., 2020; Zhang et al., 2018) calculates a scaled foot skating when in contact with the floor: $s=disp\cdot(2-2^{h/0.03})$, where $disp$ is the foot displacement in two consecutive frames, $h$ denotes the higher foot height in consecutive frames and 0.03m is the threshold value for contact. We do not calculate skate metric for sequences where the feet are not on a flat floor, such as crawling and climbing down stairs. Our method can generate the motions closest to the keyframe and show fewer skating and jerk artifacts. Our method effectively preserves the semantics specified by the text prompts, while the baseline DNO occasionally ignores the text prompts to reach the goal keyframe, as illustrated in the examples of pacing in circles and dancing in the supplementary video. This highlights the superior capability of our latent motion primitive-based DART in harmonizing spatial control and text semantic alignment.
Table 3: Quantitative evaluation of noise optimization-based In-between. The best results except dataset are in bold and ‘$\pm$’ indicates the 95% confidence interval.
[…]
Dataset - - $2.83\pm 0.00$ $0.99\pm 0.00$
OmniControl $17.22\pm 0.80$ $5.88\pm 0.57$ $5.48\pm 0.90$ $1.26\pm 0.10$
DNO $2.15\pm 0.24$ $5.52\pm 1.86$ $5.12\pm 0.55$ $0.72\pm 0.07$
DART Opt.(ours) $0\textbf{0.00}\pm 0$ 0.00 $\textbf{0.54}\pm 0.03$ $\textbf{3.97}\pm 0.96$ $\textbf{0.71}\pm 0.03$
Human-scene interaction.
We qualitatively show that our latent noise optimization control can be applied to human-scene interaction synthesis, where the goal is to control the human to interact naturally with the surrounding environment. Given the text prompts specifying the actions and durations, we use latent noise optimization to control the human to reach the goal joint location while adhering to the scene contact and collision constraints. The scenes are represented as signed distance fields for evaluating human-scene collision and contact constraints. We present generated interactions of climbing stairs and walking to sit on a chair in Fig. 3 and the supplementary video.
(a) Walk, turn left, sit on the chair
(b) Walk upstaris
(c) Walk downstaris
Figure 3: Illustrations of human-scene interaction generation given text prompts and goal pelvis joint location (visualized as a red sphere). Best viewed in the supplementary video.
4.3 Reinforcent Learning-Based Control Using DART
Combining DART with reinforcement learning-based control, we train text-conditioned goal-reaching policies for three actions including ‘walk’, ‘run’, and ‘hop on the left leg’. We evaluate the controllers on paths consisting of sequences of waypoints. We compare our method to a baseline GAMMA (Zhang & Tang, 2022), which learns a motion model using the same motion source of AMASS (Mahmood et al., 2019) dataset as ours and trains RL controllers to reach goal locations. However, GAMMA does not support text conditioning and can only generate walking motion. The evaluation metrics include the reach time, the success rate of reaching the final goal waypoint, foot skating, and foot-floor distance. The evaluation results are shown in Tab. 4. Our controller consistently reaches all goals within a reasonable timeframe, while GAMMA occasionally fails to meet the final goal and may float off the floor beyond the contact threshold. Our text-conditioned goal-reaching policy achieves a generation speed of 240 frames per second. We believe our results on text-conditioned goal-reaching controllers demonstrate the potential of DART as a foundational human motion model, upon which control models for multiple tasks can be learned without collecting separate datasets for different tasks.
Table 4: Quantitative evaluation of text-conditioned goal-reaching controller. The best results are in bold and ‘$\pm$’ indicates the 95% confidence interval.
Time (s)$\downarrow$ Success $\uparrow$ Skate (cm/s) $\downarrow$ Floor distance (cm)$\downarrow$
GAMMA walk $31.44\pm 2.58$ $0.95\pm 0.03$ $5.14\pm 1.58$ $5.55\pm 0.84$
Ours ‘walk’ $13.82\pm 0.10$ $\mathbf{1.0}\pm 0.0$ $5.07\pm 0.18$ $1.87\pm 0.02$
Ours ‘run’ $\mathbf{12.16}\pm 0.06$ $\mathbf{1.0}\pm 0.0$ $\textbf{4.70}\pm 0.13$ $2.02\pm 0.02$
Ours ‘hop on left leg’ $13.89\pm 0.12$ $\mathbf{1.0}\pm 0.0$ $41.18\pm 0.61$ $\mathbf{1.43}\pm 0.02$
5 Limitations and conclusions
Our method relies on motion sequences with frame-level aligned text annotations, as provided in BABEL, to achieve precise text-motion alignment and natural transition between actions. When trained on the coarse sentence-level motion labels from HML3D, the text-motion alignment degenerates for texts describing multiple actions, resulting in motions randomly switching between the described actions in a random order. This occurs because each short motion primitive inherently matches only a portion of the sequence’s semantics. Using a coarse sentence-level description as the text label for a primitive introduces semantic misalignment and ambiguity. Inspired by Stoffl et al. (2024), we aim to explore hierarchical latent spaces to effectively tackle both fine-grained and global sequence-level semantics in future work.
Our method DART effectively learns a text-conditioned motion primitive space that enables real-time online motion generation driven by natural languages. Additionally, the learned powerful motion primitive space allows for precise spatial motion control via latent noise optimization or reinforcement learning policies. Experiments demonstrate the superiority of DART in harmonizing spatial control with motion text-semantic alignment in the generated motions.
Appendix A Motion Primitive Representation
Representation. We represent each frame of the motion primitive as a tuple of $(\mathbf{t},\mathbf{R},\mathbf{\theta},\mathbf{J},d\mathbf{t},d\mathbf{R},d\mathbf{J})$, where $\mathbf{t}\in\mathbb{R}^{3}$ denotes the global body translation, $\mathbf{R}\in\mathbb{R}^{6}$ denotes the 6D rotation representation (Zhou et al., 2020) of the global body orientation, $\mathbf{\theta}\in\mathbb{R}^{21\times 6}$ is the 6D representation of 21 joint rotations, $\mathbf{J}\in\mathbb{R}^{22\times 3}$ denotes the 22 joints locations, $d\mathbf{t}\in\mathbb{R}^{3}$ denotes the temporal difference with previous frame’s translation, $d\mathbf{R}\in\mathbb{R}^{6}$ denotes the 6D representation of the relative rotation between current and previous frame’s body orientation, and $d\mathbf{J}\in\mathbb{R}^{22\times 3}$ denotes the temporal diffence between current and previous frame’s joint locations.
Such overparameterized motion representation has multiple benefits. Firstly, the joint rotation components $\mathbf{\theta}$ are compatible with animation pipelines, saving the time-consuming optimization-based skeleton-to-body conversion required by the commonly used HML3D (Guo et al., 2022) representation in text-to-motion methods. Moreover, including the joint location components $\mathbf{J}$ facilitates solving physical constraints like reducing foot skating and joint trajectory control. Our motion representation also models the first-order kinematics with the temporal difference features to improve the motion naturalness.
Canonicalization. We represent motion primitives in a human-centric local coordinates frame to canonicalize the primitive features and facilitate model learning. Each motion primitive is canonicalized in a local coordinates system centered at the first frame body. The origin is located at the pelvis of the first frame body, the X-axis is the horizontal projection of the vector pointing from the left hip to the right hip, and the Z-axis is pointing in the inverse gravity direction. Given the pelvis, left hip, and right hip joints, the origin is located at the pelvis joint, and the local axis system can be derived by:
Algorithm 3 Motion primitive rotation canonicalization
Input: right_hip $\in\mathbb{R}^{3}$, left_hip $\in\mathbb{R}^{3}$
x_axis = right_hip - left_hip
x_axis[2] = 0 $\triangleright$ Project to the xy plane
normalize(x_axis)
z_axis = [0, 0, 1] $\triangleright$ Inverse gravity direction
y_axis = cross_porudct(z_axis, x_axis)
normalize(y_axis)
return [x_axis, y_axis, z_axis]
We store the canonicalization transformations and apply their inverse transformation to the generated motion primitives to recover global motions in the world coordinates.
Appendix B Datasets
We train separate DART models on motion-text data from the BABEL (Punnakkal et al., 2021) and HML3D (Guo et al., 2022) dataset. Both BABEL and HML3D use motion sources from the AMASS (Mahmood et al., 2019) dataset. Their main difference is that BABEL features fine-grained frame-level text annotation while HML3D uses coarse sequence-level annotation.
The BABEL dataset contains motion capture sequences with frame-aligned text labels that annotate the fine-grained semantics of actions. Fine-grained text labels in BABEL allow models to learn precise human action controls and natural transitions among actions. We use the motion data at a framerate of 30 frames per second the same as prior works (Barquero et al., 2024; Athanasiou et al., 2022). During training, motion primitives are randomly sampled from data sequences and the text label is randomly sampled from all the action segments that overlap with the primitive. To alleviate the action imbalance in the BABEL dataset, we use the provided action labels to perform importance sampling during training so that the motion data of each action has roughly equal sampling chances despite their varying frequency in the original dataset.
The HML3D dataset contains short motions with coarse sequence-level sentence descriptions. We only use a subset of HML3D in training since its subset HumanAct12 and left-right mirroring motion sequences only provide joint locations instead of SMPL (Loper et al., 2015; Pavlakos et al., 2019) body sequences that our motion representation requires. The joint rotations used in the original HML3D are calculated using naive inverse kinematics and can not be directly used to animate human motions. We therefore only train using the subset where the SMPL body motion sequences is available. The HML3D motions are sampled at a framerate of 20 frames per second. During training, we randomly sample primitives with uniform probability and the text label is randomly chosen from one of the multiple sentence captions of the overlapping sequence.
Appendix C Motion Primitive VAE
Architecture. Our Motion Primitive VAE employs the transformer-based architecture. Both the encoder and decoder consist of 7 transformer encoder layers with skip connections (Chen et al., 2023). The transformer layers use the dropout rate of 0.1, feed-forward dimension of 1024, hidden dimension of 256, 4 attention heads, and the gelu activation function. The latent space dimension is 256. After finishing training, we follow (Rombach et al., 2022) to calculate the variance of the latent variables using a data batch. When training the latent denoiser model, the raw latent output from the encoder is scaled to have a unit standard deviation.
Losses. The motion primitive VAE is trained with the following losses:
$L_{VAE}=L_{rec}+w_{KL}\times L_{KL}+w_{aux}\times L_{aux}$ (5)
$L_{rec}=\mathcal{F}(\hat{\mathbf{X}},\mathbf{X})$ (6)
The reconstruction loss $L_{rec}$ aims to minimize the distance between the reconstructed future frames $\hat{\mathbf{X}}$ and the ground truth future frames $\mathbf{X}$, where $\mathcal{F}(\cdot,\cdot)$ denotes the distance function and we use the smoothed L1 loss(Girshick, 2015) in implementation.
$\mathcal{L}_{KL}=KL(q(z|\mathbf{H})||\mathcal{N}(0,I)),$ (7)
$L_{KL}$ penalizes the distribution difference between the predicted distribution and a standard Gaussian, where $KL(\cdot,\cdot)$ denotes the Kullback-Leibler divergence (KL), and $q(z|\mathbf{H})$ denotes the predicted distribution from the encoder $\mathcal{E}$. We use a small KL term of $1e^{-6}$ following Rombach et al. (2022) as we aim to keep the latent space expressive and only use the small KL loss to avoid arbitrarily high-variance latent spaces.
$L_{aux}=\mathcal{F}(\bar{d\mathbf{t}},\hat{d\mathbf{t}})+\mathcal{F}(\bar{d\mathbf{J}},\hat{d\mathbf{J}})+\mathcal{F}(\bar{d\mathbf{R}},\hat{d\mathbf{R}})$ (8)
The auxiliary loss $L_{aux}$ regularizes the predicted temporal difference features $\hat{d\cdot}$ of translations, global orientation, and joints to be close to the actual temporal differences $\bar{d\cdot}$ calculated from the predicted motions. We use $w_{KL}=1e^{-6}$ and $w_{aux}=100$ in our experiments. We train the motion primitive VAE with the AdamW (Kingma & Ba, 2017) optimizer and the learning rate is set to $1e^{-4}$ with linear annealing.
We conduct ablation studies on the impacts of the losses. We evaluate motion primitive VAEs trained with different loss weights on autoregressive motion sequence reconstruction error and motion jittering in reconstructed motions. Using large KL loss weight of $w_{KL}=1$, reduces the model expressiveness (the reconstruction error increases from 0.08 to 0.44 compared to $w_{KL}=1e^{-6}$ on the test motions) and fails to accurately reconstruct complex sequences like cartwheeling. Using a small KL loss weight of $w_{KL}=1e^{-6}$ can maintain the expressiveness of the learned VAE while allowing the latent distribution to deviate a bit more from a standard Gaussian. Applying the auxiliary losses with $w_{aux}=100$ helps to reduce the jittering in the reconstructions and improve the motion quality compared to using $w_{aux}=0$, as reflected by a smaller jerk metric of 2.45 when using $w_{aux}=100$ compared to a jerk of 3.67 when using $w_{aux}=0$.
Appendix D Latent Denoiser Model
D.1 Losses
We train the latent denoiser model using DDPM (Ho et al., 2020) with 10 diffusion steps and use a cosine noise scheduler. The denoiser model is trained with the following losses:
$L_{denoiser}=L_{simple}+w_{rec}\times L_{rec}+w_{aux}\times L_{aux}$ (9)
$L_{simple}=\mathcal{F}(\mathcal{G}(z_{t},t,\mathbf{H},c),z_{0}),$ (10)
where $\mathcal{F}(\cdot,\cdot)$ is a distance function and we use the smooth L1 loss (Girshick, 2015) in our implementation. We train the denoiser to predict the clean latent variable with the simple objective $L_{simple}$, and apply the feature reconstruction loss $L_{rec}$ and auxiliary losses $L_{aux}$ on the decoded motion primitive $\hat{\mathbf{X}}=\mathcal{D}(\mathcal{G}(z_{t},t,\mathbf{H},c))$ to ensure the decoded motion primitives are valid.
D.2 Scheduled Training
We use scheduled training to improve the stability of long sequence generation and the text prompt controllability. Long-term prediction stability is a significant challenge in autoregressive generation since the sample distribution can drift and accumulate during autoregressive generation. When the sample drifts out of the distribution covered by the learned model, the generation results can go wild. Our latent motion primitive model also faces the long-term stability challenge as an autoregressive method.
To alleviate the out-of-distribution problems, we use the scheduled training (Ling et al., 2020; Bengio et al., 2015; Rempe et al., 2021; Martinez et al., 2017) to progressively introduce the test-time distributions during training. Specifically, we train the latent denoiser model on sequences of $N$ consecutive motion primitives and use the prediction result of the previous motion primitive instead of the ground truth dataset to extract the history motion input $\mathbf{H}$. We name such history motion extracted from the predicted last primitive as rollout history. Using the rollout history instead of the ground truth history introduces the test-time distribution which can differ from the dataset distribution, e.g., unseen human poses or out-of-distribution combinations of human bodies and text labels. Exposing the model to such test-time distribution at training can improve the long-term generation stability and increase the text controllability when facing novel combinations of history motion and text prompts at generation time.
The scheduled training has three stages to progressively introduce the rollout history. The first stage is fully supervised training where only the ground truth history is used during training. The second scheduled learning stage randomly replaces the ground truth history motion with rollout history motion by a probability linearly increasing from 0 to 1. The third stage of rollout training always uses the rollout history instead of the ground truth history. The scheduled training algorithm for the latent denoising model is shown in Alg. 5.
Algorithm 4 Calculate rollout probability
1:Input: current iteration number $iter$, number of train iterations in the first supervised stage $I_{1}$, number of train iterations in the second scheduled stage $I_{2}$
2:Output: rollout probability $p$
3:function rollout_probability($iter$, $I_{1}$, $I_{2}$)
4:     if $iter\leq I_{1}$ then $\triangleright$ no rollout in the first supervised stage
5:         $p\leftarrow 0$
6:     else if $iter>I_{1}+I_{2}$ then $\triangleright$ the third rollout stage always use rollout
7:         $p\leftarrow 1$
8:     $\triangleright$ linearly scheduled rollout probability in the second scheduled stage
9:         $p\leftarrow\frac{iter-I_{1}}{I_{2}}$
10:     end if
11:     return $p$
12:end function
Algorithm 5 Scheduled training for the latent denoising model
1:Input: pretrained motion primitive decoder $\mathcal{D}$ and encoder $\mathcal{E}$, latent variable denoiser $\mathcal{G}_{\Theta}$ parameterized by ${\Theta}$, total diffusion steps $T$, optimizer $\mathcal{O}$, loss criterion $\mathcal{L}$, train dataset $\mathcal{X}$.
2:Scheduled training parameters: the number of train iterations in the first supervised stage $I_{1}$, the number of train iterations in the second scheduled stage $I_{2}$, the number of train iterations in the third rollout stage $I_{3}$, the maximum number of primitive rollouts $N$ during training.
3:
4:$I_{total}\leftarrow I_{1}+I_{2}+I_{3}$
5:$iter\leftarrow 0$
6:while $iter<I_{total}$ do
7:     $[\mathbf{H}_{seed},\mathbf{X}^{1},c^{1},...,\mathbf{X}^{N},c^{N}]\sim\mathcal{X}$
8:     $\triangleright$ sample $N$ consecutive motion primitives with text labels from dataset $\mathcal{X}$
9:     $\mathbf{H}\leftarrow\mathbf{H}_{seed}$triangleright initialize motion history
10:     for $i\leftarrow 1$ to $N$ do $\triangleright$ number of rollouts
11:         $z_{0}^{i}=\mathcal{E}(\mathbf{H},\mathbf{X}^{i})$triangleright compress motion primitive into latent space
12:         $t\sim\mathcal{U}[0,T)$triangleright sample diffusion step $t$
13:         $z_{t}^{i}\leftarrow\textsc{forward\_diffusion}(z_{0}^{i},t)$
14:         $\hat{z}_{0}^{i}=\mathcal{G}_{\Theta}(z_{t}^{i},t,\mathbf{H},c^{i}$) $\triangleright$ latent denoising model prediction
15:         $\hat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\hat{z}_{0}^{i})$triangleright decode predicted latent variable to future motion frames
16:         $\nabla\leftarrow\nabla_{\Theta}\mathcal{L}(z_{0}^{i},\hat{z}_{0}^{i},\mathbf{H},\mathbf{X}^{i},\hat{\mathbf{X}}^{i})$triangleright model parameter gradient calculation
17:         $\Theta\leftarrow\mathcal{O}(\Theta,\nabla)$triangleright model update using optimizer
18:
19:         $p\leftarrow$ rollout_probability($iter,I_{1},I_{2}$)
20:         $\triangleright$ update history motion using predicted or GT motion by a scheduled probability
[…]
22:              $z_{T}^{i}\leftarrow\textsc{forward\_diffusion}(z_{0}^{i},T)$triangleright maximum noising simulating inference time
23:              $\hat{z}_{0}^{i}=\textsc{ddpm\_sample}(\mathcal{G}_{\Theta},z_{T}^{i},T,\mathbf{H},c^{i})$triangleright full DDPM denoising loop
24:              $\hat{\mathbf{X}}=sg(\mathcal{D}(\mathbf{H},\hat{z}_{0}^{i}))$triangleright decode predicted latent variable and stop gradient
25:              $\mathbf{H}\leftarrow\textsc{canonicalize}(\hat{\mathbf{X}}^{F-H+1:F})$
26:         $\triangleright$ use ground truth history
27:              $\mathbf{H}\leftarrow\textsc{canonicalize}(\mathbf{X}^{F-H+1:F})$
28:         end if
29:
30:         $iter\leftarrow iter+1$
31:     end for
32:end while
We use $w_{rec}=1$ and $w_{aux}=10000$ in our experiments. The denoiser model is trained using an AdamW optimizer. The learning rate is set to $1e^{-4}$ with linear annealing. Our denoiser model is trained with scheduled training (Ling et al., 2020; Bengio et al., 2015), consisting of a fully supervised stage of 100K iterations, a scheduled stage of 100K iterations, and a rollout stage of 100K iterations. We set the maximum number of rollouts as 4. With the scheduled training, our latent motion primitive model can stably generate long motion sequences and better respond to the text prompt control even at poses that are not paired with the text prompt in the dataset.
Appendix E Text-conditioned temporal motion composition
E.1 Experiment details
We apply DART to conduct online motion generation using the rollout algorithm 1 with a default seed motion $\mathbf{H}_{seed}$ of rest standing and using a classifier-free guidance of $w=5$. We use the released checkpoints of FlowMDM (Barquero et al., 2024), TEACH (Athanasiou et al., 2022), and DoubleTake (Shafir et al., 2024) for baseline comparison. We adjust the handshake size and blending length of DoubleTake to be compatible with the shortest segment length of 15 frames. The history-conditioned modification of T2M-GPT is retrained on the BABEL dataset using the original hyperparameters. At generation time, the last frames of the previous action segment are encoded into tokens and provided as the first tokens when generating the next action segment to provide history conditioning of the previous action.
We extract the timeline of action segments for evaluation from the BABEL (Punnakkal et al., 2021) valid set. For each sequence in the validation set, we sort the original frame labels provided by BABEL to obtain a list of tuples of text prompts and durations. We randomly sample one text label when a segment is annotated by multiple texts. We skip the ‘transition’ text labels due to the ambiguous semantics and clamp the duration length with a minimum of 15 frames.
E.2 Perceptual User Studies.
We conduct user studies to quantitatively compare our method DART with baselines to provide a more comprehensive and convincing evaluation. We run binary user studies on Amazon Mechanical Turk(AMT) to evaluate generation realism and text-motion semantic alignment. Participants are given generation results from two different methods and are asked to select the generation that is perceptually more realistic or better aligns with the action text stream in subtitles, as illustrated in Fig. 4. During the realism evaluation, action descriptions were not displayed to eliminate distractions. We sample 256 sequences of action texts and durations from the BABEL dataset and use each method to generate motions given the action timelines. Participants are shown random pairs of results from our method and a baseline, and they are asked to choose the better generation. Each comparison is voted by 3 independent participants, with the video pairs randomly shuffled to ensure that participants do not know the method source.
Figure 4: Illustration of the binary user study interface for evaluating motion-text semantic alignment (top) and perceptual realism(bottom). Participants are requested to select the generation that is perceptually more realistic or better aligns with the action descriptions in subtitles (only visible in semantic user study).
E.3 Ablation Studies.
We conduct ablative studies on architecture designs, diffusion steps, primitive representation, and scheduled training on the text-conditioned motion composition task.
Removing the variational autoencoder (DART-VAE) and training diffusion model in the raw motion space results in a significantly higher jittering as reflected by the peak jerk (PJ) and Area Under Jerk (AUJ) metrics. This validates the effectiveness of integrating the variational encoder to compress the high-frequency noise in motion data and improve motion generation quality.
Without the scheduled training (DART-schedule), the model can not effectively respond to text control and have significantly worse R-Prec and FID metrics. This is because, without scheduled training, the model can easily encounter out-of-distribution combinations of history motion and text prompts during autoregressive generation.
We also include the ablation study of training a model to predict the single next frame conditioned on the current frame (per frame) as used in Shi et al. (2024). This is a special case of primitive with history length $H=1$ and future length $F=1$. The ablative model has significantly worse R-Prec and FID metrics and cannot respond to text prompts. This indicates using frame-by-frame prediction is less effective than primitives with a reasonable horizon ($H=2$, $F=8$) in learning text-conditioned motion space.
DART can learn high-quality text-conditioned motion primitive models using very few diffusion steps because of the simplicity of the primitive representation. We conduct ablative studies of training DART using different numbers of DDPM diffusion steps. Reducing the diffusion steps from 100 to fewer than 10 does not significantly affect performance. However, an extremely low diffusion step number of 2 leads to a much higher FID, indicating poorer motion quality.
Table 5: Ablation studies results on text-conditioned temporal motion composition. The first row includes the metrics of the dataset for reference. Symbol $\rightarrow$’ denotes that closer to the dataset reference is better and ‘$\pm$’ indicates the 95% confidence interval.
Segment Transition
[…]
Dataset 0.00$\pm$0.00 0.72$\pm$0.00 8.42$\pm$0.15 3.36$\pm$0.00 0.00$\pm$0.00 6.20$\pm$0.06 0.02$\pm$0.00 0.00$\pm$0.00
Ours 3.79$\pm$0.06 0.62$\pm$0.01 8.05$\pm$0.10 5.27$\pm$0.01 1.86$\pm$0.05 6.70$\pm$0.03 0.06$\pm$0.00 0.21$\pm$0.00
DART-VAE 4.23$\pm$0.02 0.62$\pm$0.01 8.33$\pm$0.12 5.29$\pm$0.01 1.79$\pm$0.02 6.73$\pm$0.23 0.20$\pm$0.00 0.96$\pm$0.00
DART-schedule 8.08$\pm$0.09 0.39$\pm$0.01 8.05$\pm$0.12 6.96$\pm$0.03 7.41$\pm$0.10 6.58$\pm$0.06 0.03$\pm$0.00 0.18$\pm$0.00
per frame($H$=1,$F$=1) 10.31$\pm$0.09 0.29$\pm$0.01 6.82$\pm$0.13 7.41$\pm$0.01 7.82$\pm$0.09 6.03$\pm$0.08 0.02$\pm$0.00 0.08$\pm$0.00
steps 2 4.44$\pm$0.04 0.60$\pm$0.00 8.20$\pm$0.15 5.38$\pm$0.01 2.24$\pm$0.02 6.77$\pm$0.07 0.05$\pm$0.00 0.20$\pm$0.00
steps 5 3.49$\pm$0.09 0.63$\pm$0.00 8.25$\pm$0.15 5.18$\pm$0.01 2.11$\pm$0.07 6.74$\pm$0.11 0.05$\pm$0.00 0.20$\pm$0.00
steps 8 3.70$\pm$0.03 0.62$\pm$0.01 8.04$\pm$0.13 5.25$\pm$0.03 2.15$\pm$0.08 6.72$\pm$0.15 0.06$\pm$0.00 0.20$\pm$0.00
steps 10 (Ours) 3.79$\pm$0.06 0.62$\pm$0.01 8.05$\pm$0.10 5.27$\pm$0.01 1.86$\pm$0.05 6.70$\pm$0.03 0.06$\pm$0.00 0.21$\pm$0.00
steps 50 3.82$\pm$0.05 0.60$\pm$0.00 7.74$\pm$0.07 5.30$\pm$0.01 2.11$\pm$0.10 6.58$\pm$0.10 0.06$\pm$0.00 0.22$\pm$0.00
steps 100 4.16$\pm$0.06 0.61$\pm$0.00 7.82$\pm$0.15 5.32$\pm$0.02 2.20$\pm$0.05 6.43$\pm$0.10 0.06$\pm$0.00 0.21$\pm$0.00
Appendix F Latent diffusion noise optimization-based control
Optimization details. We use the Adam (Kingma & Ba, 2017) optimizer with a learning rate of 0.05 to optimize latent noises $Z_{T}$ for 100 to 300 steps. The learning rate is linearly annealed to 0 and the gradient is normalized to stabilize optimization. The optimization running time is dependent on both the motion sequence length and the number of optimization steps. An example experiment of a target motion sequence of 60 frames and 100 optimization steps costs around 74 seconds.
Human-scene interaction. In human-scene interaction synthesis, we use two scene constraints to avoid human-scene collision and foot-floating artifacts. To reduce human-scene interpenetration, $cons\_coll(\cdot)$ penalizes joints colliding into scenes with negative SDF value and can be formulated as:
$cons\_coll(\mathbf{M})=-\sum_{J}(\Psi(J(\mathbf{M}))-J(\mathbf{M})_{thresh})_{-}$ (11)
where $\Psi(\cdot)$ is the signed distance function returning the signed distance from the query point to the closest scene point. $J(\cdot)$ extracts the joint locations from the motion $\mathbf{M}$. $(\cdot)_{thresh}$ is the joint-dependent contact threshold value, which is determined by the joint-skin distance in the rest pose. $(\cdot)_{-}$ denotes clipping positive values.
To reduce the occurrence of foot-floating artifacts, $cons\_cont(\cdot)$ encourages the foot to be in contact with the scene and is defined as:
$cons\_cont(\mathbf{M})=\sum_{F}(\Psi(F(\mathbf{M}))-F(\mathbf{M})_{thresh})_{+}$ (12)
where $F(\cdot)$ extracts the lower foot joint location. $(\cdot)_{+}$ denotes clipping negative values.
Appendix G Reinforcement Learning-based Control
Architecture. Both the actor and critic networks are 4-layer MLPs with residual connections and a hidden dimension of 512. We apply the tanh scaling: $x=4\cdot tanh(x)$ (Ling et al., 2020) to the actor output to clip the action prediction in the range of $[-4,4]$, avoiding unbounded action predictions. The actor networks are initialized with close to zero weights to boost policy training following (Andrychowicz et al., 2020). The observation input contains a 512D text embedding, a 552D of history motion, a 1D observation of floor height relative to the human pelvis, a 1D observation of the floor plane distance from the pelvis to the goal location, and a 3D unit vector of the goal direction in the human-centric coordinates frames. We clamp the goal distance observation with a maximum value of 5m and the goal direction in the range of a 120-degree field of view to simulate egocentric human perception.
Rewards. The text-conditioned goal-reaching control policies are trained to maximize the discounted expectation of the following rewards (Zhang & Tang, 2022; Li et al., 2024; Zhao et al., 2023):
We use three distance-related rewards to encourage the human agent to minimize its distance to the goal location.
$r_{dist}=D^{i}-D^{i-1}$ (13)
The distance reward $r_{dist}$ encourages the human to get closer to the goal location, where $D^{i}$ is the 2D distance between the human pelvis and the goal location at step $i$.
$r_{succ}=\begin{cases}1&\text{if }D^{i}<0.3\\ 0&\text{otherwise}\end{cases}$ (14)
The success reward $succ$ gives a sparse but strong reward when the human arrives at the goal, where 0.3m is the success threshold value.
$r_{ori}=\frac{\langle p^{i}-p^{i-1},g-p^{i-1}\rangle+1}{2},$ (15)
The moving orientation reward encourages the moving orientation to be aligned with the goal orientation, where $p^{i}$ is the human pelvis location at step $i$, and $g$ is the goal location.
Moreover, we apply scene constraints-related rewards to discourage unnatural behaviors such as foot skating and collision with the floor.
$r_{skate}=-disp\cdot(2-2^{h/0.03}),$ (16)
The skate reward $r_{skate}$ penalizes foot displacements when in contact with the floor, where $disp$ is the foot displacement in two consecutive frames, $h$ denotes the higher foot height in two consecutive frames and 0.03m is the threshold value for contact.
$r_{floor}=-(|lf|-0.03)_{+},$ (17)
The floor contact reward $r_{floor}$ penalizes when the lower foot distance to the floor is above the threshold of 0.03m, where $lf$ denotes the height of the lower foot, ||$\cdot$|| ⋅ | denotes the absolute value operator, and $(\cdot)_{+}$ is the clipping operator with a minimum of 0.
The rewards are weighted with $w_{dist}=1$, $w_{succ}=1$, $w_{ori}=0.1$, $w_{floor}=100$, $w_{skate}=10$ for the hopping controller and 100 for the others.
Appendix H Computing resources
Our experiments and performance profiling are conducted on a workstation with single RTX 4090 GPU, intel i7-13700K CPU, 64GiB memory. The workstation runs with Ubuntu 22.04.4 LTS system.